OpenStudy (anonymous):

how do you find something raised to the -1 power?

OpenStudy (anonymous):

You take the mulitplicative inverse of the expression. For example: a^-1 = 1/a

OpenStudy (anonymous):

ok so 2 raised to -1 is 1/2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks i compeltely forgot

OpenStudy (anonymous):

And 2 raised to -2 is 1/4.

OpenStudy (anonymous):

that i remember but i just forgot -1

OpenStudy (anonymous):

wait but when does something raised to something power equal 1?

OpenStudy (anonymous):

raised to the 0 power.

OpenStudy (anonymous):

thanks it's all coming back to me now

OpenStudy (anonymous):

Negative exponents are inverses (x^-1 = 1/x). Fractional exponents are roots (x^(1/2) is sqrt(x)). 0 exponents are 1 (x^0 = 1).

OpenStudy (anonymous):

And we would go on with the rules: (a)^b*(a^c) = a^(b+c) (a^b)^c = a^(bc) Important rules, even though they seem very innocent.

OpenStudy (agreene):

Phi, you can also use what most post WWII number theorists use... We define exponents in the following logarithmic method: \[\large n^x=e^{x \ln (n)}\] putting in 0 for x: \[n^0 = e^{0 \ln(n)}\] \[n^0 = e^{0}\] \[\therefore n^0=1\space |\space \forall : n \ne 0 \]

OpenStudy (phi):

That will mean anything to a kid in high school?

OpenStudy (agreene):

Well, in order to understand recursion... one must first understand recursion... in order to understand recursion... It's a rather straightforward collapsing proof... so I'd like to think they'd get it.

OpenStudy (agreene):

And maybe it will help them remember the rules of exponents, cause in that general logarithmic form... it self defines every exponential form, which is fancy.